OpenStudy (anonymous):

cot theta = 1. Where theta lies in Quadrant 3. What is csc feta?

OpenStudy (anonymous):

-root(2)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[Cot \theta=1\] \[Tan \theta=1\] \[Sin \theta/Cos \theta=1\] \[Sin \theta=Cos \theta\] \[Cosec \theta=1/Sin \theta\] \[Cosec \theta=-1\] In the third quadrant only Tan/Cot functions are positive rest everything is negative.

OpenStudy (anonymous):

Btw I am sure you mean Cosec Theta and not Feta? Lol just curious

OpenStudy (anonymous):

oopps, I had a little typo, but yea I meant that

OpenStudy (anonymous):

Haha yeah well I was wondering what's feta xD anyways it's -1 I think because sin/cosec functions in third quadrant are -ve.

OpenStudy (anonymous):

kk thanks